OpenStudy (anonymous):

A ball is thrown vertically upward from the top of a 100-foot tower, with an initial velocity of 20 ft/sec. Its position function is s(t) = –16t2 + 20t + 100. What is its velocity in ft/sec when t = 1 second? –12 –44 100 –32

OpenStudy (cwrw238):

first find the derivative ds/dt which gives the the formula for the velocity then plug in t = 1

OpenStudy (cwrw238):

can you find the derivative/

OpenStudy (cwrw238):

have you learnt some calculus ? differentiation?

OpenStudy (anonymous):

it is NOT B -32

OpenStudy (anonymous):

the correct answer is A) -12 using ( f(a+h)-f(a) ) / h, plug in a = 1 and solve >> >> ( h(-32 - 16h +20) ) / h cancel and plug in h = 0 -12